METCRAX: Logbook Entries

METCRAX: Radiometers Messages: 29 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
194 Tue 02-Sep-2008Net Rad Testing Continuedaster
191 Thu 17-Apr-2008New Nrads out at MarshallSemmer
189 Wed 12-Mar-2008Post Q7 testsAllSemmer
185 Tue 15-May-2007Horizontal Rnet at wl had wrong coefficientsWest_LowerSemmer
183 Mon 23-Apr-2007Radiometer maintenance by SebastianAllHoch
182 Mon 09-Apr-2007Licor shadowband sensorsAllSemmer
180 Wed 27-Dec-2006Rlw has been quality controlled excluding the rfd tower on floorAllaster
178 Tue 26-Dec-2006Tsfc.pyrg has been quality controlledAllaster
176 Tue 26-Dec-2006Rdiff mostly quality controlledAllaster
175 Tue 26-Dec-2006Rsw has been quality controlledAllaster
174 Tue 26-Dec-2006Rnet is mostly quality controlledAllaster
162 Wed 01-Nov-2006RFD testRFDOncley
154 Mon 30-Oct-2006radiometer [post posting]Horst
152 Sun 29-Oct-2006rad measurementsAllOncley
148 Sat 28-Oct-2006Rnet slope anglesAllSemmer
138 Thu 26-Oct-2006logger change at ELEast_LowerSemmer
136 Wed 25-Oct-2006el logger died againEast_LowerOncley
123 Tue 24-Oct-2006logger wires brokeEast_UpperOncley
121 Tue 24-Oct-2006Battery change at logger ELEast_LowerSemmer
85 Mon 16-Oct-2006Fixed RFD configurationFloorHorst
59 Thu 12-Oct-2006Fixed Tcase.out at elEast_LowerHorst
53 Wed 11-Oct-2006configuration corrected for RFD tower at flr; dsm restartedFloorMartin
47 Sun 08-Oct-2006Plugged in Tsurface at swSouthwestHorst
22 Fri 06-Oct-2006radiometers cleanedRFDHoch
21 Fri 06-Oct-2006radiometers cleanedFloorHoch
20 Fri 06-Oct-2006radiometers cleanedEast_UpperHoch
19 Fri 06-Oct-2006radiometers cleanedEast_LowerHoch
14 Thu 05-Oct-2006Rnet and Tsurface at swSouthwestHorst
3 Tue 03-Oct-2006Installed net radiometerSouthwestHorst

194: Radiometers, Site , Tue 02-Sep-2008 11:59:19 MDT, Net Rad Testing Continued
August 29, 2008

3:14 - 3:51pm MDT

Chris fixed Net Rad. testing site.  Noticed that wrong Net Rads.  One of the Q7s is part of ISS'.  Will not be able to test that one this week.  I will come back to the site and do a re-test on WL.
191: Radiometers, Site , Thu 17-Apr-2008 17:10:47 MDT, New Nrads out at Marshall
April 18, 2008

2:40 - 4:18 MDT

Chris installed new Net Rads to testing platform.  Now testing East Lowers.

Q97035 (East) and Q90076 (West/Horz.).

New coeff. installed.

189: Radiometers, Site All, Wed 12-Mar-2008 07:41:07 MDT, Post Q7 tests
Dave Whiteman asked if we could do some follow-up tests on the Q7 sensors
because data analysis of the horizontal Q7s deployed on the slopes is
Out at Marshall, a 4 component radiation system has been deployed along with
three Q7 units. The 4 component system along with the center Q7 is the same
system deployed at the floor site of METCRAX. The 2 other Q7s are from west
upper. The west upper sensors are evenly spaced between the 4-comp and the
end of the dark horse.
The plan is to rotate in the other Q7 units over the next 3 to 4 weeks
depending on sky conditions.

Test 1:  East-Q94196, WestQ97034, CentralQ99258

185: Radiometers, Site West_Lower, Tue 15-May-2007 15:02:51 MDT, Horizontal Rnet at wl had wrong coefficients
I decided to use west lower as the radiation system for CHATS and did a check 
of the coefficients. I notice an error in the values used for the horizontal Q7.They had coefficients of 8.57 (+) and 10.53(-). The values should have been 
9 (+) and 11.22 (-). I am assuming that the Q7 locations were not changed 
during setup. 

183: Radiometers, Site All, Mon 23-Apr-2007 15:07:08 MDT, Radiometer maintenance by Sebastian
Here are Sebastian's collected notes:

radiometer cleaning:

10/07/2006 15:21-15:31 cleaned RFD tower instruments

10/07/2006 15:27-15:29 cleaned 4-component radiometers floor, not net

10/08/2006 06:51 cleaned 4-component radiometers floor, not net,
condensation on sw_in.

10/10/2006 14:48 cleaned 4-component radiometers floor, not net

10/15/2006 10:55 adjusted levelling of net radiometer sw

10/18/2006 15:08 cleaned RFD tower instruments. Re-adjusted level of 0.5 m
boom, other booms OK. Checked shadowbands rim and floor.

10/26/2006 16:05 cleaned 4-component radiometers floor.

10/26/2006 16:10 cleaned RFD tower instruments on 0.5 m and 2 m booms.

10/28/2006 13:42-1354 cleaned RFD tower instruments.

Shadowband mainentance:

10/11/2006 13:15 adjusted shadowband rim (Barringer point). Licor sensor
1/3 exposed to direct.

10/11/2006 13:42 adjusted shadowband floor, very thin edge of Licor
exposed to direct.

10/15/2006 11:53 checked shadowband rim. slight correction, Licor not
exposed to direct.

10/15/2006 15:43 checked shadowband floor. slight correction, Licor not
exposed to direct, re-adjusted levelling of Licor, levelled net.

10/18/2006 15:08 Checked shadowbands rim and floor.

10/28/2006 13:15 adjusted shadowband rim (Barringer point), no direct on

10/28/2006 13:40 adjusted shadowband floor, no direct on Licor.

182: Radiometers, Site All, Mon 09-Apr-2007 14:08:40 MDT, Licor shadowband sensors
During METCRAX 2 Licor pyranometers were used in a shadowband mode at the rim
site and the floor site. These sensors were sent back to Licor for
calibrations after the project. Below are the results of those calibrations.
During the calibration Licor noticed a scratch on the diffuser of PY7122.They
went ahead and repaired the sensor before any calibration. This resulted in
a major calibration change for this sensor. Therefore, the pre-project
calibration should be used for this sensor.

sensor     PRE        POST
PY7122    -134.97     -104.34    (new diffuser for post)
PY108335  -103.88     -104.34

180: Radiometers, Site All, Wed 27-Dec-2006 20:31:14 MST, Rlw has been quality controlled excluding the rfd tower on floor
The Rlw has been quality controlled excluding the rfd tower on floor

1. had a bad Rpile for a little bit and had a little data removed when the sensor was disconnected, everything else ok.
2. Rlw.out: Rlw.out.el had data removed because of a bad Tcase. Rlw.out.flr had a random outlier removed and had some bad data removed when the sensor was disconnected. Everything else ok.
178: Radiometers, Site All, Tue 26-Dec-2006 18:37:29 MST, Tsfc.pyrg has been quality controlled
Tsfc.pyrg has been quality controlled

1. Everything is ok, only removed one outlier at Tsfc.pyrg.el
176: Radiometers, Site All, Tue 26-Dec-2006 18:14:15 MST, Rdiff mostly quality controlled
Most of Rdiff is quality controlled

1. When sebastian adjusted sensors on 10/11 at ~13:00, readings changed 
2. Removed a spike from Rdiff.flr
3. Everything else ok
175: Radiometers, Site All, Tue 26-Dec-2006 17:16:19 MST, Rsw has been quality controlled
The short wave radiation has been quality controlled

1., some outliers removed, but everything else ok
2. Rsw.out, some outliers removed for Rsw.out.flr, but everything else ok

174: Radiometers, Site All, Tue 26-Dec-2006 16:14:06 MST, Rnet is mostly quality controlled
The Rnet is mostly quality controlled except for changing the cal coefficients
on the sw (and wl.hor) site.

1. Rnet.flr had little data removed because of an outlier and releveling, 
2. had some bad data when the sensor was disconnected
3. Rnet.sw, in addition to the bad cal coefficient, there was some bad data 
before the sensor was hooked up, releveling, and flipping over to read cal 
4. Everything else ok

TWH, 5/15/07:  See logbook entry 185, Rnet.hor.wl also had wrong coefficients. 

162: Radiometers, Site RFD, Wed 01-Nov-2006 16:33:58 MST, RFD test
About 1100 yesterday, we removed the upper RFD sensors.  Later that day,
Sebastian moved all of them to the same height (~1.5m) for an intercomparison
experiment.  This will end when we shut down flr tomorrow.

154: Radiometers, Site , Mon 30-Oct-2006 11:09:53 MST, radiometer [post posting]
On September 28, Dave Whiteman requested the re-orientation of  five
sawhorses -- floor, lower and upper east, and lower and upper west. The
remaining two were not adjusted -- the rim and the sodar site. The floor
sawhorse was moved to the northeast to reduce reflections from the salt
pan that was to the south of its original position. Adjustments to the
slope sawhorses were made as follows. First, we oriented the sawhorse
rails to an azimuth angle parallel to the fall line of the crater
sidewalls above the site, considering the fall line on a large scale
that was representative of the crater as a whole (i.e, ignoring any
local topography under the sawhorse). Secondly, we adjusted the
inclination angle of the rail so that it was parallel to the general
slope of the sidewall at that elevation, ignoring any local wrinkles in
the underlying topography. Thus, we think that the radiometers on the
sawhorses are now oriented so that they are parallel to the underlying
large-scale slope of the crater sidewalls.

We shot the azimuth and inclination angles of the slope site sawhorses
and got some GPS readings, as follows:

site, azimuth, inclination, lat, long, elevation
upper west, 249, 24
lower west, 232, 07
lower east, 257, 08, 35deg 01.629min, 111deg 01.190min, 5132 ft
upper west, 249, 24
Barringer Pt, ?, 00?, 35deg 01.784min, 111deg 01.750min, 5688 ft

152: Radiometers, Site All, Sun 29-Oct-2006 12:46:46 MST, rad measurements
Yesterday, I tried to make all the measurements needed for the radiometers.
Hopefully, this is enough!

(Note that pitch for rad is roll for beam and vice versa.)
(Note that the GPS elevations do not always make sense.)

5/7/07, TWH: Note that the datascope was set to a declination of 12.2 deg E,
whereas the NOAA web site states that the declination should be 11.4 deg E

Net Rad (horiz): pitch 0deg, roll 0deg
Beam: pitch 0deg, roll 1deg, azimuth 88deg
Rad position: 35d 01.785', 111d 01.751', 1740m elev [GPS accuracy 6m]
Station position: 35d 01.767', 111d 01.747', 1741m elev [5m]

Net Rad (tilted): pitch 1d (up), roll 4d
Beam: pitch 4d, roll 0d, az 69.7d
Rad: 35d 01.637', 111d 01.530', 1574 elev [5m]
Station: 35d 01.648', 111d 01.530', 1576m elev [?]

Net Rad (tilted): pitch 1d (up), roll 23d
Beam: pitch 21d, roll 0d, azimuth 88.3d
Rad: 35d 01.645', 111d 01.622', 1608m elev [9m]
Station: 35d 01.646', 111d 01.612, 1610m elev [11m!]

Net Rad (tilted): pitch 0d, roll 23d, az 354.0d
Beam: pitch 25d, roll 0d
Rad: 35d 01.633', 111d 01.105', 1606m elev [5m]
Station: 35d 01.638', 111d 01.105', 1598m elev [6m]

Net Rad (tilted): pitch 2d (up), r 8d, az 356.0
Beam: pitch 6d, roll 1d
Rad: 35d 01.631', 111d 01.188', 1573m elev [6m]
Station: 35d 01.637', 111d 01.185', 1573m elev [6m]

Rad: 35d 01.683', 111d 01.348', 1565m elev [5m]
Station: 35d 01.681', 111d 01.360', 1586m elev [5m]
RFD: 35d 01.680', 111d 01.356', 1562m elev [5m]

148: Radiometers, Site All, Sat 28-Oct-2006 09:58:35 MST, Rnet slope angles
A measurement of the slope angle for the rnets at eu, el, wl, and wu was done.
A second set of measurments will be done again.
site	slope (degrees)
wu	21
wl	5
el	5
eu	22.5
138: Radiometers, Site East_Lower, Thu 26-Oct-2006 14:22:47 MST, logger change at EL
The logger at EL now has a direct link to the ADAM. So far no drop out.
This occurred at ~12:45 local today.
136: Radiometers, Site East_Lower, Wed 25-Oct-2006 16:27:31 MST, el logger died again
died at about 1245 and still not back up.  The hard-wired logger does respond
to +++ commands, so the problem probably is on the logger side.  (Though the
logger battery was 13.2V and rising when it died.)

We're a bit puzzled.

123: Radiometers, Site East_Upper, Tue 24-Oct-2006 15:26:51 MST, logger wires broke
While turning off the logger for RF testing, the power wires broke at the 
switch.  I shoved them back in, but this wiring needs to be redone.

121: Radiometers, Site East_Lower, Tue 24-Oct-2006 14:38:02 MST, Battery change at logger EL
Steve and I decided to change the logger battery at EL to see if it is causing
the logger outages during the middle of the day. This took place during a
crater visit between 11:30 and 13:00 local time.

85: Radiometers, Site Floor, Mon 16-Oct-2006 11:37:28 MST, Fixed RFD configuration
Charlie corrected the aster and flr configurations to properly parse 
the RFD radiometer data.

59: Radiometers, Site East_Lower, Thu 12-Oct-2006 17:21:37 MST, Fixed Tcase.out at el
Found a loose wire in the el Campbell data logger which was coming from Pyg.out.
Screwed it down and fixed the Tcase.out
53: Radiometers, Site Floor, Wed 11-Oct-2006 12:35:02 MST, configuration corrected for RFD tower at flr; dsm restarted
The configuration had an incorrect scanfFormat for
the data logger from the RFD tower, which has caused
the Tpile to be put into the Rpile variable, and the
Tpile.8.5m to be skipped entirely. 

Fixed the configuration, downloaded to flr, and
restarted the dsm.

47: Radiometers, Site Southwest, Sun 08-Oct-2006 15:41:24 MST, Plugged in Tsurface at sw
Tsurface at sw was not plugged in!  Rectified this oversight.
22: Radiometers, Site RFD, Fri 06-Oct-2006 11:01:29 MST, radiometers cleaned
03.10. 16:48 Cleaned RDF tower instruments
21: Radiometers, Site Floor, Fri 06-Oct-2006 10:53:46 MST, radiometers cleaned
03.10. 16:18 cleaned 4 comp radiometers, CG4s , not net rad, noticed 4 comps to be somewhat loose. 
20: Radiometers, Site East_Upper, Fri 06-Oct-2006 10:50:35 MST, radiometers cleaned
03.10. 15:21 cleaned 4-component radiometers, not net radiometers

19: Radiometers, Site East_Lower, Fri 06-Oct-2006 10:46:44 MST, radiometers cleaned
03.10 15:15 cleaned 4 components, not net radiometers
14: Radiometers, Site Southwest, Thu 05-Oct-2006 15:55:00 MST, Rnet and Tsurface at sw
Around 1500 local time, I turned the Rnet over to read the coefficients:
positive = 8.70
negative = 10.77

Around 1530 local time, I removed power from the logger to add power to
the Everest Tsurface plug.  Plugged in Tsurface.
3: Radiometers, Site Southwest, Tue 03-Oct-2006 17:51:51 MST, Installed net radiometer
Around 19:00 local time, I installed a net radiometer at the southwest site.
Note that I did not enter the coefficient in the logger.

5/10/07, TWH:  Note that the time, 19:00, has to be wrong because the entry is
dated 17:51:51. My guess from looking at the Rnet data is that this should be
around 17:00.